# Training Convolutional Neural Networks with the Forward-Forward Algorithm

### 25 Dec 2023 v2

@@ Appendix B Contribution of the First Layer
Figure 9: Our implementation of FF-trained CNNs does not require the inclusion of the goodness of the first layer during training. Continuous lines represent evolution of the discrimination accuracy during the training phase, when the first layer is not included. Dashed lines represent the discrimination accuracy evolution if its goodness is included.
@@ Appendix C Effect of Max Pooling
We also examine how max pooling layers influence FF-trained CNNs by testing five different configurations. Two exhibit an increased architectonic complexity by having the number of filters increase for increasing layer number. Our analysis, shown in Table 2, indicates a decrease in accuracy when max pooling layers were applied, suggesting that preserving the entire information is preferable.
@@
Table 3: Adding max pooling layers reduces the performance of FF trained CNN. Five sets of hyperparameters were trained using either only three convolutional layers or by interleaving each pair of convolutional layers with a max pooling layer. Unlike the rest of the paper, the last two configurations have different numbers of convolutions in the three layers. Batch size is 50 and a linear classifier is used for inference.

### 07 Jan 2024 v3

@@ Appendix B Contribution of the first layer
Figure 9: Our implementation of FF trained CNNs does not require the inclusion of the goodness of the first layer during training. Continuous lines represent evolution of the discrimination accuracy during the training phase, when the first layer is not included. Dashed lines represent the discrimination accuracy evolution if its goodness is included.
@@ Appendix C Effect of max pooling
We also examine how max pooling layers influence FF trained CNNs by testing five different configurations and increased architectonic complexity by having higher number of filters for increasing layer number. As shown in Table 3, our analysis exhibits a decrease in accuracy when max pooling layers were applied, suggesting that preserving the entire information is preferable.
@@
Table 3: Adding max pooling layers reduces the performance of FF trained CNN. Five sets of hyperparameters were trained using either only three convolutional layers, or by interleaving each pair of convolutional layers with a max pooling layer. Unlike the rest of the paper, the last two configurations have different numbers of convolutions in the three layers. Batch size is 50 and a linear classifier is used for inference. Five different sets of hyperparameters were trained, with the best results highlighted in bold.